A friend thought of two numbers, one of which is 8 greater than the other. When he trembled the smaller number, it was 4 larger than the greater number. What are the two numbers my friend thought of?
Answer by addingup(3677)

one number: x
the other: x+8
When you triple (I think that's what you mean, not tremble)
3x = x+8+4
3x = x+12
2x = 12
x = 6 This is the small number, and the other, the problem says:
6+8 = 14
`````````````````````````````````````````````````
check:
6*3 = 18
and
18-14 = 4 correct.
J